# Orbital orders under magnetic fields in cubic PrIr2Zn20

### v3

@@ Appendix A Multiple-q configuration
In this Appendix, we show the real-space modulations for the double-$\bm{q}$ and quadruple-$\bm{q}$ orders since they are rather complicated. The real-space configurations for the double- and quadrupole-$\bm{q}$ orders are given by Eq. (25). Here, we concentrate on the $A$ sublattice moments, since those for the $B$ sublattice are obtained by ${\boldsymbol{\phi}}_{n}\to{\boldsymbol{\phi}}_{n}(-1)^{p_{n}}$. Thus, one can rewrite Eq. (25) as
@@
Here, we have assumed, without loss of generality, double-$\bm{q}$ orders with ${\bm{k}}_{1}$ and ${\bm{k}}_{2}$. Since $\bm{r}$ is the position in the fcc lattice, there are four choices for $\mathbb{E}_{2}(\bm{r}):(\pm 1,\pm 1)$, and eight choices for $\mathbb{E}_{4}(\bm{r}):\pm(1,1,1,1),\ \pm(1,1,-1,-1),\ \pm(1,-1,1,-1)$, and ±$\pm(-1,1,1,-1)$. In Fig. 9, $\mathbb{E}_{\mathcal{N}}(\bm{r})$’s are illustrated and the colors distinguish the vectors $\mathbb{E}_{2,4}$. Note that only the $A$-sublattice positions in the diamond structure are shown. The unit cell for the double-$\bm{q}$ orders contains four $A$ sublattice sites, each of which forms ferroic chain along the direction parallel to ${\bm{k}}_{3}+{\bm{k}}_{4}\parallel(0,-1,1)$ as shown in Fig. 9(a). The primitive translation vector along this direction is unchanged, while those perpendicular to it are modified to $\frac{1}{2}(2,1,1)$ and $\frac{1}{2}(2,-1,-1)$. For quadruple-$\bm{q}$ orders, the primitive translational vectors are as twice large as those for the original fcc lattice: (1,1,0), (1,0,1), and (0,1,1). The unit cell contains eight sites for the quadruple-$\bm{q}$ orders and sixteen when the $B$-sublattice sites are included.
Figure 9: (a) Double-$\bm{q}$ configuration $\mathbb{E}_{2}$ at the $A$ sublattice. Tubes connect the same $\mathbb{E}_{2}$’s. Thin lines indicates the NN bonds of diamond structure. Arrows indicates the direction of ${\bm{k}}_{1}$ and $-{\bm{k}}_{2}$. (b) Quadruple-$\bm{q}$ configuration $\mathbb{E}_{4}$. (c) Symmetric quadruple-$\bm{q}$ configuration for $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,0,0)$ and (d) for $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,\pi/2,3\pi/2)$. For (d), there are four kinds of ferroic chains. Two of them are similar to the double-$\bm{q}$ configuration for $x=2n$($n:$ integers) planes with the $2z^{2}-x^{2}-y^{2}$ orbital, while the other two chains for $x=2n+1$ plane with the $x^{2}-y^{2}$ orbital run along ${\bm{k}}_{1}+{\bm{k}}_{2}$ direction.
Since several double-$\bm{q}$ configurations are shown in Fig. 8, we here examine the real space configuration of the symmetric quadruple-$\bm{q}$ orders discussed in Sec. IV.1, although they are not stabilized in the $T$$H$ phase diagram in this paper. It is evident that the configurations with “collinear” types such as $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,0,0)$ and $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,0,\pi)$ are partially ordered states. For example, the quadrupole moments vanish at the sites where $\sum_{n}e^{i{\bm{k}}_{n}\cdot\bm{r}}=0$ for the former as shown in Fig. 9(c). As an example of “non-collinear” case, Fig. 9(d) illustrates the configuration for $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,\pi/2,3\pi/2)$.
@@ Appendix B Renormalization effects due to the fields at the X points
The free energy for the X points fields are sublattice diagonal and $F_{\rm X}=\sum_{\alpha=A,B}F_{\rm X}^{\alpha}$, where
$F^{\alpha}_{\rm X}=$ $\frac{1}{2}\sum_{\ell={\rm X,Y,Z}}\bm{Q}^{\alpha}_{\ell}\begin{bmatrix}a_{\rm X}-6bu_{0}^{\alpha}&6bv_{0}^{\alpha}\\ 6bv_{0}^{\alpha}&a_{\rm X}+6bu_{0}^{\alpha}\end{bmatrix}\bm{Q}^{\alpha}_{\ell}$
@@
Here, the fourth-order terms and the cubic term for three X points fields are neglected since the modes at the X points are not critical. $a_{\rm X}$ and $C$’s and $S$’s consisting of two L points variables are given as
@@
where $(\ell,m,n)=({\rm X},1,2),\ ({\rm Y},1,4),\ ({\rm Z},1,3)$, and $(\ell,m^{\prime},n^{\prime})=({\rm X},3,4),\ ({\rm Y},2,3),\ ({\rm Z},2,4)$. By minimizing $F_{\rm X}$ with respect to the X points fields, we obtain their stationary values as
@@
Substituting them into Eq. (32), we obtain up to the linear order in $\phi_{0}$: $\delta F=\sum_{\alpha={\rm A,B}}\delta F^{\alpha}_{0}+\delta F^{\alpha}_{1}+\cdots$ as
$\delta F^{\alpha}_{0}=$ $-\frac{18b^{2}}{a_{\rm X}}\sum_{m>n}(\phi^{\alpha}_{m}\phi^{\alpha}_{n})^{2}$
@@
$\delta F_{1}^{\alpha}=$ $-\frac{108b^{3}}{a_{\rm X}^{2}}\Bigg{[}\sum_{m>n}\cos(2\theta^{\alpha}_{mn}-\theta_{0}^{\alpha})(\phi^{\alpha}_{m}\phi^{\alpha}_{n})^{2}\phi_{0}^{\alpha}$
$+6\cos(\bar{\theta}^{\alpha}-\bar{\theta}_{\alpha})\phi^{\alpha}_{1}\phi^{\alpha}_{2}\phi^{\alpha}_{3}\phi^{\alpha}_{4}\phi_{0}^{\alpha}\Bigg{]},$ (37)
@@
The terms appearing in $\delta F_{1}^{\alpha}$ are emergent fifth-order coupling constants between the mode at the $\Gamma$ and the L points. Since we have kept terms up to the fourth orders in the original free energy in Eq. (6), cases for large $108b^{3}/a_{\rm X}$ are beyond the applicable range of the approximation.
@@ Appendix C Free energy for the orthogonal double-q configuration
We show the free energy for the orthogonal double-$\bm{q}$ orders with $\bm{\phi}_{2}\perp\bm{\phi}_{1}$: $\bm{\phi}_{1}=\phi_{1}(\cos\theta,\sin\theta),\bm{\phi}_{2}=\phi_{2}(\cos(\theta+\tfrac{\pi}{2}),\sin(\theta+\tfrac{\pi}{2}))=\phi_{2}(-\sin\theta,\cos\theta)$, with the $\Gamma$ point mode $\bm{\phi}_{0}=\phi_{0}(\cos\theta_{0},\sin\theta_{0})$. Here, we assume $\bm{\phi}_{1}=\bm{\phi}_{1}^{A}=\bm{\phi}_{1}^{B}$ and $\bm{\phi}_{2}=\bm{\phi}_{2}^{A}=-\bm{\phi}_{2}^{B}$ as discussed in the main text. The free energy is given as
$F=$ $\sum_{n=1,2}\Big{[}(a_{\rm L}+J_{\rm L}+\lambda_{c}^{(1)}\phi_{0}^{2})\phi_{n}^{2}+\frac{c}{2}\phi_{n}^{4}\Big{]}+\frac{c^{\prime}}{2}\phi_{1}^{2}\phi_{2}^{2}$
@@
$+\frac{216b^{3}}{a_{\rm X}^{2}}\phi_{0}\phi^{2}_{1}\phi^{2}_{2}\cos(4\theta-\theta_{0})+F_{\Gamma},$ (39)
where $c^{\prime}=2c-72b^{2}/a_{\rm X},\ \lambda_{b}=b,\ \lambda_{c}^{(1)}=\lambda_{c}^{(2)}=c$, $F_{\Gamma}$ is given by Eq. (10). Let us briefly discuss possible stationary angle $\theta$ under magnetic fields. For simplicity, we assume $\theta_{0}=0$, since $\tilde{h}$ induces $u_{0}$ under both $\bm{H}\parallel[001]$ and $[110]$. We treat $\phi_{0,1,2}$ as given parameters and the stationary condition of $F$ with respect to $\cos(2\theta)$ leads to
@@
$F_{{\rm III}_{2q}}=$ $2\big{[}a_{\rm L}+J_{\rm L}+(\lambda_{c}^{(1)}+\lambda_{c}^{(2)})\phi_{0}^{2}\big{]}\phi_{1}^{2}$
$+\left(c+\frac{c^{\prime}}{2}-\frac{216b^{3}}{a_{\rm X}^{2}}\phi_{0}\cos\theta_{0}\right)\phi_{1}^{4}+F_{\Gamma}.$ (41)
Although we will not discuss in detail, we can also see that the deviation $\delta\phi=\phi_{1}-\phi_{2}\neq 0$ from the phase III2q accompanies changes in the angle as $\theta=\pi/4+\epsilon$. The quadratic form of $\delta\phi$ and $\epsilon$ describe the stability of the phase III2q and when one of the eigenvalues vanishes a second-order transition takes place toward the phase II2q.
@@
$F_{\rm{I}_{2q}}=$ $\sum_{n=1,2}\Big{[}(a_{\rm L}+J_{\rm L}+\lambda_{c}^{(1)}\phi_{0}^{2})\phi_{n}^{2}+\frac{c}{2}\phi_{n}^{4}\Big{]}+\frac{c^{\prime}}{2}\phi_{1}^{2}\phi_{2}^{2}$
@@
$+\frac{216b^{3}}{a_{\rm X}^{2}}\phi_{0}\phi^{2}_{1}\phi^{2}_{2}\cos\theta_{0}+F_{\Gamma}.$ (42)

### v4

@@ Appendix A Multiple-q configuration
In this Appendix, we show the real-space modulations for the double-$\bm{q}$ and quadruple-$\bm{q}$ orders since they are rather complicated. The real-space configurations for the double- and quadrupole-$\bm{q}$ orders are given by Eq. (25). Here, we concentrate on the $A$-sublattice moments, since those for the $B$ sublattice are obtained by ${\bm{\phi}}_{n}\to{\bm{\phi}}_{n}(-1)^{p_{n}}$. Thus, one can rewrite Eq. (25) as
@@
Here, we have assumed, without loss of generality, double-$\bm{q}$ orders with ${\bm{k}}_{1}$ and ${\bm{k}}_{2}$. Since $\bm{r}$ is the position in the fcc lattice, there are four choices for $\mathbb{E}_{2}(\bm{r}):(\pm 1,\pm 1)$, and eight choices for $\mathbb{E}_{4}(\bm{r}):\pm(1,1,1,1),\ \pm(1,1,-1,-1),\ \pm(1,-1,1,-1)$, and $\pm(-1,1,1,-1)$. In Fig. 9, $\mathbb{E}_{\mathcal{N}}(\bm{r})$’s are illustrated and the colors distinguish the vectors $\mathbb{E}_{2,4}$. Note that only the $A$-sublattice positions in the diamond structure are shown in Fig. 9(b). The unit cell for the double-$\bm{q}$ orders contains four $A$-sublattice sites, each of which forms ferroic chain along the direction parallel to ${\bm{k}}_{3}+{\bm{k}}_{4}\parallel(0,-1,1)$ as shown in Fig. 9(a). The primitive translation vector along this direction is unchanged, while those perpendicular to it are modified to $\frac{1}{2}(2,1,1)$ and $\frac{1}{2}(2,-1,-1)$. For quadruple-$\bm{q}$ orders, the primitive translation vectors are as twice large as those for the original fcc lattice: (1,1,0), (1,0,1), and (0,1,1). The unit cell contains eight sites for the quadruple-$\bm{q}$ orders and sixteen when the $B$-sublattice sites are included.
Since several double-$\bm{q}$ configurations are shown in Fig. 7, we here examine the real space configuration of the symmetric quadruple-$\bm{q}$ orders discussed in Sec. IV.1, although they are not stabilized in the TTHH phase diagram in this paper. It is evident that the configurations with “collinear” types such as $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,0,0)$ and $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,0,\pi)$ are partially ordered states. For example, the quadrupole moments vanish at the sites where $\sum_{n}e^{i{\bm{k}}_{n}\cdot\bm{r}}=0$ for the former as shown in Fig. 9(c); $\bm{\phi}^{A}(\bm{r})={\bf 0}$ for four fcc sites and $\bm{\phi}^{B}(\bm{r})=-\tfrac{1}{2}\bm{\phi}^{A}(\bm{0})$. As an example of “non-collinear” case, Fig. 9(d) illustrates the configuration for $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,\pi/2,3\pi/2)$. This is also a partially ordered state, where $\bm{\phi}^{B}(\bm{r})=\bm{0}$ at two of the $B$ sublattice sites within the cubic unit cell shown in Fig. 9(d), while the other two have finite moments $|\bm{\phi}^{B}(\bm{r})|=\sqrt{2}|\bm{\phi}^{A}(\bm{0})|$.
@@ Appendix B Renormalization effects due to the fields at the X points
The free energy for the X-points fields are sublattice diagonal and ${\mathcal{F}}_{\rm X}=\sum_{\alpha=A,B}{\mathcal{F}}_{\rm X}^{\alpha}$, where
${\mathcal{F}}^{\alpha}_{\rm X}=$ $\frac{1}{2}\sum_{\ell={\rm X,Y,Z}}\bm{Q}^{\alpha}_{\ell}\begin{bmatrix}a_{\rm X}-6bu_{0}^{\alpha}&6bv_{0}^{\alpha}\\ 6bv_{0}^{\alpha}&a_{\rm X}+6bu_{0}^{\alpha}\end{bmatrix}\bm{Q}^{\alpha}_{\ell}$
@@
Here, the fourth-order terms and the cubic term for three X-point fields are neglected since the modes at the X points are not critical. $a_{\rm X}$ and $C$’s and $S$’s consisting of two L-point variables are given as
@@
where $(\ell,m,n)=({\rm X},1,2),\ ({\rm Y},1,4),\ ({\rm Z},1,3)$, and $(\ell,m^{\prime},n^{\prime})=({\rm X},3,4),\ ({\rm Y},2,3),\ ({\rm Z},2,4)$. By minimizing ${\mathcal{F}}_{\rm X}$ with respect to the X-point fields, we obtain their stationary values as
@@
Substituting them into Eq. (32), we obtain up to the linear order in $\phi_{0}$: $\delta{\mathcal{F}}_{\rm{XL}}=\sum_{\alpha={\rm A,B}}\delta{\mathcal{F}}^{\alpha}_{\rm{XL}0}+\delta{\mathcal{F}}^{\alpha}_{\rm{XL}1}+\cdots$ as
$\delta{\mathcal{F}}^{\alpha}_{\rm{XL}0}=$ $-\frac{18b^{2}}{a_{\rm X}}\sum_{m>n}(\phi^{\alpha}_{m}\phi^{\alpha}_{n})^{2}$
@@
$\delta{\mathcal{F}}_{\rm{XL}1}^{\alpha}=$ $-\frac{108b^{3}}{a_{\rm X}^{2}}\Bigg[\sum_{m>n}\cos(2\theta^{\alpha}_{mn}-\theta_{0}^{\alpha})(\phi^{\alpha}_{m}\phi^{\alpha}_{n})^{2}\phi_{0}^{\alpha}$
$+6\cos(\bar{\theta}^{\alpha}-\theta_{0}^{\alpha})\phi^{\alpha}_{1}\phi^{\alpha}_{2}\phi^{\alpha}_{3}\phi^{\alpha}_{4}\phi_{0}^{\alpha}\Bigg],$ (37)
@@
The terms appearing in $\delta{\mathcal{F}}_{\rm{XL}1}^{\alpha}$ are emergent fifth-order coupling constants between the mode at the $\Gamma$ and the L points. Since we have kept terms up to the fourth orders in the original free energy in Eq. (6), cases for large $108b^{3}/a_{\rm X}$ are beyond the applicable range of the approximation. We also note that the last terms in Eqs. (36) and (37) vanish when the sublattice ($\alpha$) summations are carried outwith assuming $\bm{\phi}_{n}^{B}=(-1)^{p_{n}}\bm{\phi}_{n}^{A}$. See the discussions in Secs. III.1 and IV.1.
@@ Appendix C Free energy for the orthogonal double-q configuration
We show the free energy for the orthogonal double-$\bm{q}$ orders with $\bm{\phi}_{2}\perp\bm{\phi}_{1}$: $\bm{\phi}_{1}=\phi_{1}(\cos\theta,\sin\theta),\bm{\phi}_{2}=\phi_{2}(\cos(\theta+\tfrac{\pi}{2}),\sin(\theta+\tfrac{\pi}{2}))=\phi_{2}(-\sin\theta,\cos\theta)$, with the $\Gamma$-point mode $\bm{\phi}_{0}=\phi_{0}(\cos\theta_{0},\sin\theta_{0})$. Here, we assume $\bm{\phi}_{1}=\bm{\phi}_{1}^{A}=\bm{\phi}_{1}^{B}$ and $\bm{\phi}_{2}=\bm{\phi}_{2}^{A}=-\bm{\phi}_{2}^{B}$ as discussed in the main text. The free energy is given as
${\mathcal{F}}_{\rm or}=$ $\sum_{n=1,2}\Big[(a_{\rm L}+J_{\rm L}+\lambda_{c}^{(1)}\phi_{0}^{2})\phi_{n}^{2}+\frac{c}{2}\phi_{n}^{4}\Big]+\frac{c^{\prime}}{2}\phi_{1}^{2}\phi_{2}^{2}$
@@
$+\frac{216b^{3}}{a_{\rm X}^{2}}\phi_{0}\phi^{2}_{1}\phi^{2}_{2}\cos(4\theta-\theta_{0})+{\mathcal{F}}_{\Gamma},$ (39)
where $c^{\prime}=2c-72b^{2}/a_{\rm X},\ \lambda_{b}=b,\ \lambda_{c}^{(1)}=\lambda_{c}^{(2)}=c$, and ${\mathcal{F}}_{\Gamma}$ is given by Eq. (10). Let us briefly discuss possible stationary angle $\theta$ under magnetic fields. For simplicity, we assume $\theta_{0}=0$, since $\tilde{h}$ induces $u_{0}$ under both $\bm{H}\parallel[001]$ and $[110]$. We treat $\phi_{0,1,2}$ as given parameters and the stationary condition of ${\mathcal{F}}_{\rm or}$ with respect to $\cos(2\theta)$ leads to
@@
${\mathcal{F}}_{{\rm III}_{2q}}=$ $2\big[a_{\rm L}+J_{\rm L}+(\lambda_{c}^{(1)}+\lambda_{c}^{(2)})\phi_{0}^{2}\big]\phi_{1}^{2}$
$+\left(c+\frac{c^{\prime}}{2}-\frac{216b^{3}}{a_{\rm X}^{2}}\phi_{0}\cos\theta_{0}\right)\phi_{1}^{4}+{\mathcal{F}}_{\Gamma}.$ (41)
Although we do not discuss this in detail, we can also see that the deviation $\delta\phi=\phi_{1}-\phi_{2}\neq 0$ from the phase III2q accompanies changes in the angle as $\theta=\pi/4+\epsilon$. The quadratic form of $\delta\phi$ and $\epsilon$ describes the stability of the phase III2q. When one of the eigenvalues vanishes, a second-order transition takes place toward the phase II2q.
@@
${\mathcal{F}}_{{\rm{I}}_{2q}}=$ $\sum_{n=1,2}\Big[(a_{\rm L}+J_{\rm L}+\lambda_{c}^{(1)}\phi_{0}^{2})\phi_{n}^{2}+\frac{c}{2}\phi_{n}^{4}\Big]+\frac{c^{\prime}}{2}\phi_{1}^{2}\phi_{2}^{2}$
@@
$+\frac{216b^{3}}{a_{\rm X}^{2}}\phi_{0}\phi^{2}_{1}\phi^{2}_{2}\cos\theta_{0}+{\mathcal{F}}_{\Gamma}.$ (42)